If 15% of the population carries at least one copy of the recessive allele: What is the predicted frequency of individuals in the population that express the dominant phenotype? What is the predicted frequency of individuals in the population that express the recessive phenotype?
1
Expert's answer
2020-04-21T12:27:19-0400

According to Hardy-Weinberg equations:

p + q = 1

p2 + 2pq + q2 = 1

where p - frequency of the dominant allele, q - frequency of the recessive allele, p2 - frequency of homozygous dominant genotype, 2pq - frequency of heterozygous genotype, q2 - frequency of recessive genotype.

As 15% of the population carries a copy of the recessive allele (q=0.15), 100% - 15% = 85% of the population carries a copy of the dominant allele (p=0.85).

As a result, the predicted frequency of individuals in the population that express the dominant phenotype equals p2 + 2pq = 0.852 + 2 × 0.85 × 0.15 = 0.9775 = 0.98

The predicted frequency of individuals in the population that express the recessive phenotype equals q2 = 0.152 = 0.0225 = 0.02.


Answer: 0.98 and 0.02